### v5

# Building Better: Avoiding Pitfalls in Developing Language Resources when Data is Scarce

Abstract
Language is a symbolic capital that affects people’s lives in many ways Bourdieu (1977, 1991). It is a powerful communication means that accounts for identities, cultures, traditions, and societies in general. Hence, data in a given language should be viewed as more than a collection of tokens. Good data collection and labeling practices are key to building more human-centered and socially aware technologies. While there has been a rising interest in under-resourced languages within the NLP community, work in this space has to overcome unique challenges such as data scarcity and access to suitable annotators. In this paper, we collect feedback from those directly involved in and impacted by NLP artefacts for medium and low-resource languages. We conduct a quantitative and qualitative analysis of the responses and highlight the main issues related to (1) data quality such as linguistic and cultural data suitability; and (2) the ethics of common annotation practices such as the misuse of online community services. Based on these findings, we make several recommendations for the creation of high-quality language artefacts that reflect the cultural milieu of its speakers, while simultaneously respecting the dignity and labor of data workers.
1 Introduction
Figure 1: The main stakeholders and their role in a data annotation project. We reached out to the CL community in our survey. It is designed for NLP researchers and practitioners who have worked on low- to medium-resource languages. Some of the questions focus on the perspectives of the subset highlighted in the figure, i.e., speakers who focus on their own languages.
Over the past few years, there has been an increasing interest in improving the scope of NLP research with more human-centered design choices Kotnis et al. (2022) and social awareness Yang et al. (2024). As language technologies depend on the quality of the data Hirschberg and Manning (2015) and their alignment with the needs of the speakers, researchers, and other users, the perspectives of these different stakeholders are key to high-quality tools and resources. That is, data selection, collection, annotation, and design choices are traditionally made by the researchers who develop the different artefacts. However, the involvement of those whose native languages are in question is paramount to better design practices Bird and Yibarbuk (2024) since language is part of their culture and identity Bourdieu (1991). That said, when dealing with low to mid-resource languages in NLP, researchers often make use of the datasets available without necessarily looking into their quality or questioning their suitability, mainly due to resource scarcity. Although progress in NLP for English and other high-resource languages has led to improving standards for corpora quality control and research practices Gebru et al. (2021); Bender and Friedman (2018); Mohammad (2022), one cannot claim the same about the data sources and prevailing practices for low- to mid-resource languages given the current research scope in the field Joshi et al. (2020). Therefore, the NLP artefacts developed for under-resourced languages and underrepresented cultures often suffer from a lack of social considerations and cause over-generalisations due to the over-reliance on data and tools that fail to incorporate the predominant linguistic and cultural features of a given language Bender and Friedman (2018), which may hinder critical progress. This can further lead to inequality Blasi et al. (2022); Held et al. (2023), sub-optimal experiences with language technologies, and could reinforce a legacy of language hierarchy Kahane (1986).
In this position paper, we shed light on the current limitations of NLP research for low- to mid-resource languages in terms of appropriate data collection, ethical annotation practices, and overall data quality. We reached out to the NLP community involved in NLP projects on under-served languages and conducted a survey to report on the common incentives, limitations, applied norms, and practices (see Figure 1). We outline the survey and present its results. Finally, based on the survey responses, we provide a set of recommendations that focus on (1) fairness and centering of the speakers of the language, (2) choosing suitable data sources, (3) setting fair and realistic expectations when recruiting annotators, and (4) avoiding cultural misrepresentation.
2 Related Work
Work on ethical practices in AI, ML, and NLP research covers a variety of topics, such as artefact documentation Bender (2011); Bender and Friedman (2018); Gebru et al. (2021); Rogers et al. (2021); Mohammad (2022) and recommendations for best practices Hollenstein et al. (2020); Mohammad (2023). Those that focus on low-resource languages are centered on the general state of NLP research in the area Held et al. (2023); Joshi et al. (2020); Blasi et al. (2022); Doğruöz and Sitaram (2022), limitations in specific tasks such as machine translation Mager et al. (2023), LLM research Mihalcea et al. (2024), or on the essential question of including people whose languages are in question Mager et al. (2023); Bird (2020, 2022); Bird and Yibarbuk (2024); Lent et al. (2022). Such work sheds light on the peculiarities of low-resource languages with the majority being vernacular languages rather than institutionalised or written Bird and Yibarbuk (2024); Bird (2024). They further advocate for language communities to take over their languages Schwartz (2022); Markl et al. (2024); Mihalcea et al. (2024). For instance, Bird and Yibarbuk (2024) focus on how experts (e.g., linguists, computer scientists) interact with the language communities using participatory design approaches Winschiers-Theophilus et al. (2010), and Cooper et al. (2024) provide recommendations on how to engage with indigenous communities without merely focusing on accuracy. Doğruöz and Sitaram (2022) further point out the importance of not treating language technologies for low-resource languages as scaled-down versions of high-resource ones and account for their specificity. Similarly, Adebara and Abdul-Mageed (2022) make related claims with a focus on features that are peculiar to African languages (e.g., tones). In addition to the language speakers, other work focuses on users such as Blaschke et al. (2024) who highlight the needs of dialect speakers and the importance of involving them in designing language technologies. Moreover, Yang et al. (2024) define social awareness and advocate for refraining from treating language in NLP as a computational problem only. In this paper, we strengthen the above discussion by shifting the focus to the practical challenges faced by NLP researchers and practitioners working on low and mid-resource languages by borrowing practices from social science Cetina (1999) to study the methodological practices and issues in the field. To the best of our knowledge, there is limited work investigating NLP research for low-resource languages while trying to connect to online communities, except for three case studies of participatory frameworks (e.g., Masakhane), and work by Lent et al. (2022), who analyse 38 responses by Creole speakers on their experiences with language technologies. By analysing the respondents’ feedback, we present practical recommendations that emphasise transparency and ethically grounded practices for building more human-centered NLP artefacts for under-served languages.
3 Survey
Our main goal is to investigate the current issues and problematic practices in NLP research for low- to mid-resource languages and provide potential solutions. Therefore, we reached out to the NLP community (i.e., *CL networks) from June to October 2024 on X, LinkedIn, Google groups, and Slack channels of online NLP communities, and by direct emails. We targeted researchers and annotators working on low- and mid-resource languages, language variants, dialects, and vernaculars, and surveyed how research is conducted. Participants report on common practices, incentives, and issues that stand out. Then, we present a quantitative and qualitative analysis of the responses.
3.1 Respondents
Respondents are NLP researchers and practitioners involved in the data collection, annotation, model construction, or other research questions related to under-served languages.
3.2 Survey Structure
We ask the respondents about (1) their previous experiences in the area, (2) current problems and limitations related to their language(s) of interest, (3) the motivation behind their involvement in various projects, and (4) how they were credited for tasks that are often specific to low-resource languages, e.g., compensation for annotations done via online community forums or participatory frameworks.
Note that we left it to the respondents to decide on what represents a mid- to low-resource language since there is no standard definition. E.g., most researchers would agree that Tamasheq is a low-resource language, while they may not have the same opinion about Malaysian.
3.2.1 General Questions
Respondents could optionally fill in their names and contact information for a potential follow-up. Then, they were asked about:
  • the language(s) they work on,
  • the project(s) they were involved in,
  • whether they are/were part of any online community (i.e., participatory research frameworks),
  • whether the project(s) they worked on are from industry, academia, or both,
  • the kind of NLP tools that are or would be relevant and useful in their language(s) of interest,
  • the reason(s) why they work on this/these language(s).
3.2.2 Reporting on Incentives and Potential Limitations
We investigate the common reasons why researchers work on low-resource languages. Therefore, we ask the participants to report on:
  • the incentive(s) for working on their language(s) of interest,
  • the incentive(s) for working on specific projects.
As we are aware of potential drawbacks in NLP for mid- to low-resource languages Blasi et al. (2022), we examined whether the respondents work in the area due to any limitations observed in available NLP tools in their language(s) of interest. Note that these questions were optional as researchers may work on low- to mid-resource languages for various other reasons. We asked the participants to report on:
  • any observed limitations and optionally list some tools or resources in their language(s) of interest as examples,
  • potential language-specific challenges in their language(s) of interest.
3.2.3 Reporting on Credit Attribution
We asked the respondents about how often they were properly credited for their work. Further, as reaching out to online participatory frameworks is common to projects that focus on under-resourced languages, we asked whether the participants were involved in past projects through participatory frameworks. This is because involving communities in NLP and ML projects is relatively new to the field and can therefore be abused (purposefully or not) due to the lack of clear standards regarding data workers in such contexts. Therefore, our questions were the following:
  • How often did the respondents receive credit for their contributions? E.g., financial compensation for annotating data.
  • How often were they offered authorship when making substantial contributions to the data collection and/or data annotation?
  • What were their incentives for projects in which they did not receive financial compensation or authorship?
  • How long did the process take especially when they were not properly credited?
4 Findings
Projects in Task Motivation
Industry 12% Data creation 47% Scientific interest 81%
Academia 57% Data annotation 33% Building language technologies 72%
Both 31% Data collection 33% Limitations in language(s) of interest 60%
Model construction 9% LLM research 59%
Table 1: Reported project affiliations, tasks in which the annotators were involved, and the different motivations or incentives. Note that percentages do not sum up to one as respondents could report on more than one project in both industry and academia.
We received 81 responses from researchers working on a wide range of languages and language families. Even though including contact information was optional, more than 90% of the respondents chose not to reply anonymously, and 80% asked for updates on the project. Table 1 shows the distribution of responses to questions on project affiliations, the tasks in which the respondents were involved, and their motivations for working on under-resourced languages. Note that percentages do not sum up to one as respondents could report on more than one project. That is, participants could be involved in several tasks and projects. As shown in Table 1, most participants were involved in dataset curation mainly motivated by scientific interest or curiosity, and for building language technologies because of observed limitations in resources dedicated to their language(s) of interest.
Figure 2: The main locations where the languages of interest of our respondents are spoken.
4.1 General Information
4.1.1 Projects
The respondents could report on one or many projects they have been involved in. As shown in Table 1, Most respondents have worked on academic projects, with a third on collaborations between industry and academia or both types of projects. [^1] Note that although ¿50% of the respondents named the projects they participated in, we do not disclose it to protect the anonymity of our respondents.
4.1.2 Languages
Among the 81 responses, the respondents worked on >70 low-resource languages they specifically named (see Appendix). Figure 2 illustrates the main locations where these languages are spoken. These include variants, dialects, and vernaculars (e.g., country-specific Arabic dialects), truly low-resource languages (e.g., Welsh, Yoreme Nokki, Setswana), and low- to medium-resource ones (e.g., Amharic, Indonesian). In addition, about 12% of the respondents reported working on language families and language branches such as South Asian languages, all Gaeilige dialects, or Arabic/English variations. A high percentage of the respondents work on high-resource languages in parallel.
4.2 Incentives and Potential Limitations
When asking the respondents about why they work on NLP for under-resourced languages, we provide them with a checklist from which they could choose more than one option or add their own entry. We report on the frequent motivations and practices that are only adopted in low- to mid-resource settings due to, e.g., data scarcity. We identify problematic instances and analyse the possible reasons behind some.
Figure 3: Frequency of each incentive. Note that the percentages do not sum up to 100 as the respondents could choose more than one option.
When further examining the common motivations, we report more detailed numbers in Figure 3. Among those who were motivated by scientific curiosity or interest in Table 1 there were those whose interest was in NLP/CL/ML research (68%) and those whose interest was in languages (68%). Note that the two are not mutually exclusive.
Moreover, for the respondents whose motivation was building language technologies, most of them were more interested in building technologies for their own language(s) (60%) as opposed to building technologies for as many languages as possible (38%). This is particularly interesting as it constitutes evidence of the power of language as a symbolic capital Bourdieu (1991), which can sometimes manifest in the feeling of “a duty” that one has towards their language. Other frequent motivations include marked limitations in language resources and tools in the language(s) of interest (60%) and the willingness to contribute to LLM research (59%).
4.2.1 Reported Limitations
More than 60% of the respondents reported working on low-resource languages due to marked limitations in currently available resources for their language(s) of interest. To shed light on these limitations, we showed the respondents a predefined list of possible shortcomings as well as a text box where they could add any observed limitations. As shown in Figure 4: the predominant limitation is data scarcity (78%); followed by the lack of representativeness of the data (58%) that can manifest in, e.g., unnatural or translationese data instances; the under-performance of the available tools (54%); their misalignment with the users’ needs (54%); the low quality of the annotations (25%); and the lack of the usefulness of the data (18%).
Figure 4: Frequency of each reported limitation when the respondents reported working on NLP for low-resource languages due to marked shortcomings.
4.2.2 Qualitative Analysis of the Limitations
Figure 5: Respondents on getting credit for projects they were involved in.
Respondents on getting credit for projects they were involved in. We provided the respondents with free text sections where they could report examples of tools or resources that suffer from the limitations that they mentioned to justify their choices. When manually processing the answers, we noticed the following recurring topics:
  1. 1.
    Limitations related to the currently available resources: such as their unavailability, small size, limited representativeness, and quality.
  2. 2.
    Limitations related to the practices adopted when building new resources: such as:
    • the reliance on machine translation tools and LLMs to build resources;
    • the lack of awareness of culture-specific and linguistic challenges of the languages in question;
    • the challenges with annotator recruitment due to the lack of availability of native or near-native speakers on commonly used annotation platforms (e.g., AMT and Prolific),
    • the potential misuse of online communities and participatory framework services.
  3. 3.
    Fundamental problems related to NLP research on low- to mid-resource languages: such as the lack of funding often due to the “low prestige” language dilemma—the false notion that some languages or language varieties are more important than others. Interestingly, this was equally observed in projects from academia and industry.
We discuss all three of these themes below.
Currently Available Resources
Since many under-resourced languages are not institutional but rather vernacular Bird and Yibarbuk (2024), collecting data presents considerable challenges when one solely relies on textual data for, e.g., Bantu languages. Further, the focus on English and the reliance on translated data harms the quality of the generated datasets as they do not capture the subtle peculiarities of a given language. It is important to note that what is translated and whether it was further verified by a fluent speaker makes a difference as translating Wikipedia texts can be easier than translating conversational, informal, or religious texts Hutchinson (2024).
Limitations with respect to Building New Resources
Lack of representativeness and unnaturalness of the data were commonly reported in the responses. The respondents reported a lack of awareness of language variants and cultural aspects when building a language-specific artefact; the reliance on the standardised version of a given language due to power dynamics (more power in the hands of well-funded institutions and established researchers); the presence of offensive utterances in the data due to a lack of data filtering; and potentially wrong assumptions about a language or a culture. For instance, the time-specific context and usage of some languages, such as ancestral ones (e.g., Coptic), have considerably changed and one has to take these facts into account. In addition, datasets may be collected from inadequate sources or could be aligned with Western values, standards, or expectations. This can be due to power differentials or a lack of deeper examination carried along with locals and native speakers. Finally, researchers rely on personal connections as it is hard to impossible to find fluent speakers of low- to mid-resource languages on commonly used annotation platforms such as Amazon Mechanical Turk and Polific. Added to this reason, the lack of funding leads researchers to turn to participatory frameworks. This practice has been at the center of major NLP contributions in recent years Birhane et al. (2022). However, despite its benefits for people with common research interests, the absence of well-established standards puts vulnerable community members at risk as their efforts may not be properly recognised.
Fundamental Problems
Many respondents reported that conducting research in low- to mid-resource languages often entails high costs of data curation and potential outreach to local communities. Further, when resources for an under-resourced language exist, they are often not freely available.
4.3 Credit Attribution
We asked the respondents to share whether they were properly credited for their work by, e.g., getting financial compensation for a long annotation task, getting involved in the writing of a research paper for a resource that they built, etc. As shown in Figure 5, most respondents (>67%) report this not being the case at least once. Figure 6 shows the distributions of responses pertaining to how the respondents were incentivised to perform an annotation task for which they were eventually not given due credit.
Problematic Incentivisation
Figure 6: Respondents on incentives when no proper credit (e.g., financial compensation for data annotation) was offered. We show the counts of the incentives and the time it took participants to complete their work for a given project (from <=2 hours to more than a month).
For the respondents who reported that they did not receive proper credit for projects in which they were involved, we report the incentives for joining these projects and the time it took the participants to complete the work. As shown in Figure 6, they were either:
  1. 1.
    a member of a community or a participatory framework (see paragraph below),
  2. 2.
    acknowledged on the website or the research paper, or
  3. 3.
    somehow manipulated into thinking that there was a professional benefit in joining without proper compensation as we explain below.
The Issue with the Over-reliance on Online Communities
When using standard crowdsourcing platforms such as AMT or Prolific, one can operationalise the annotation for a given task. Despite their shortcomings Fort et al. (2011); Irani (2015), one can attempt to protect workers by using tests and training when annotating hard tasks. However, for mid- to low-resource languages, platforms such as AMT and Prolific often do not have enough speakers registered on the platform. Therefore, researchers opt for personal connections or participatory efforts instead. On the one hand, personal outreach and community efforts make people feel more included and trust can be built more easily. On the other hand, there is a high risk of exploitation and emotional manipulation in such a case. For instance, some junior researchers reported cases of exploitation due to them being told that joining an online community that builds language resources is prestigious and worth adding to their CVs. We note that some respondents shared their frustration in the responses as they were offered, e.g., small company gifts for months of labor. As shown in Figure 6, 40% of the respondents, who spent 1 day to more than a month annotating data report negative experiences. That is, their work was not properly compensated, acknowledged, or recognised. This calls for a need to set guidelines and standards when using community or participatory research services.
5 Recommendations
While there has been a considerable amount of work on the ethics of best practices for building NLP and ML artefacts Bender and Friedman (2018); Leech et al. (2024); Mohammad (2022, 2023), our findings substantiate the fact that research on under-resourced languages presents additional challenges linked to the reliance on unconventional practices.
5.1 Center the People
Our findings show that various issues ought to be addressed early as research in the area lacks established standards and is subject to power differentials (e.g., some researchers reported their work being used by more notable individuals for fame). Many under-resourced languages are from what is called “the Global South” with a large number of them being spoken rather than written.
Speakers
Language is an important part of a population’s identity and technologies dealing with it have a direct impact on people’s lives. Past NLP work highlights how to engage with speakers and communities whose languages are in question Bird (2020); Bird and Yibarbuk (2024); Bird (2024); Cooper et al. (2024); Ramponi (2024). We further reinforce this argument with our findings, i.e., most researchers researchers report marked limitations in their languages (see Figure 4). Hence, when a researcher reaches out to a group of people with little background knowledge of their culture or language, one needs to approach these problems from the perspective of the community in question Bird (2022). The question of who is exactly served needs to be addressed early on to avoid any misconception of perceived needs for language technologies.
Researchers vs. Data Workers
In addition to the large percentage of our survey respondents who reported not being properly credited for their labor (Figure 5), there were cases of emotional manipulation (e.g., making emotional arguments such as how one’s labor will help the speakers of the language and that is compensation enough). One has to set rules and expectations with clear communication on the purpose of a given research project. For instance, when dealing with online participatory research communities for data collection and annotation, extra care needs to be shown and benevolent prejudice such as depicting oneself as a savior of a local community Bird (2022) must be avoided. Companies and research labs relying on communities for annotation and data creation need to properly compensate the contributors by, e.g., drawing legal contracts. The question of who is annotating what has to be addressed as well. The scarcity of annotators can lead to poor choices as very often, native speakers cannot be found online easily which has led to researchers choosing people from associated regions—people who do not necessarily speak the language variant in question. This results in a problematic overgeneralisation that puts different languages under the same umbrella simply because they have one or a small set of attributes in common. This often results in potentially over-simplistic solutions. For instance, variations of Arabic differ considerably but numerous research projects have treated entire regions, such as North Africa, as a monolith (e.g., to appear to have more data).
5.2 Be Fair: Give Credit where Credit is Due
Our findings show the unfortunate trend of data workers and NLP researchers suffering from a lack of recognition, especially those who are part of online communities (Figure 3). A needed follow-up work would be extensive fieldwork with the various online communities. Hence, our recommendation is a call for action on the setup of fair and comprehensive practices in participatory research projects, while taking power differentials into account. That is, existing authorship standardsneed to be followed and discussed prior to the start of a project as to whether a data worker should be listed as an author. This is particularly critical for junior researchers who substantially contribute to the resource construction and may assume that they will have a leading role in the paper. Moreover, proper financial compensation needs to be provided for annotators. Ideally, a resource paper should provide proof that the annotators were paid and treated fairly as recommended by Rogers et al. (2021).
5.3 Choose the Jargon Carefully and Be Aware of False Generalisations
As previously discussed in 5.1, it is important to embrace social awareness and avoid grouping people from colonial and Western perspectives Bird (2020, 2022); Held et al. (2023). Hence, one can avoid dismissive and outdated terms and classifications, e.g., “the rest of the world”. Note also that The World’s Values Survey classification Haerpfer and Kizilova (2012), which is often used in NLP papers (e.g., Santy et al. (2023)), presents an orientalist view of the world Said (1977). It has clear flaws such as including Christian-majority countries (e.g., Ethiopia, Rwanda) in a so-called "African-Islamic" category and groups of countries that have little in common in one category (e.g., Kyrgyzstan and Tunisia), leading to misrepresentations and enforcing stereotyping.
5.4 Set Fair and Realistic Expectations
As pointed out by Doğruöz and Sitaram (2022), tools for low-resource languages are often perceived as scaled-down versions of high-resource ones. Adding to previous work elaborating on what this may mean to the speakers Bird (2022); Markl et al. (2024), we focus on the impact of setting these expectations for researchers and practitioners. That is, they may be expected to build models similar to those built for high-resource languages, i.e., tackling the same NLP tasks, and performing extremely well. However, this can be unrealistic for various reasons such as the users’ needs Blaschke et al. (2024), the language’s specific features Bird and Yibarbuk (2024), and the lack of funding linked to the “prestige” of the language as reported by our respondents and similarly discussed by Mihalcea et al. (2024) in the context of LLM research.
No Prescription
Joshi et al. (2020) conducted a survey on the state of NLP in various languages. They found that people do not necessarily want the tools that researchers think they need. However, we should not be limiting what NLP research on low- to mid-resource languages should be about. The most challenging part is to find a balance between what technologies local communities need and the research explored by the NLP community. This balance could be reinforced by continued communication with local communities Bird and Yibarbuk (2024); Lent et al. (2022); Mager et al. (2023); Cooper et al. (2024).
Dealing with a "Solved" Problem in a New Language is an Actual Contribution
Just because automatic systems obtain high scores on some English datasets (often making use of massive amounts of English data) does not mean that the problem is solved generally, or that working on that problem in other languages is no longer interesting or valuable. Working with each new language presents many new challenges, e.g., a rich morphology or the presence of tones Adebara and Abdul-Mageed (2022). Therefore, such work should not be undervalued by calling it “a replication” study.
5.5 Check the Source Even if the Language is Low-resource
Due to the limited amount of online data available for some languages, there is a tendency to use any available online source to build resources for these languages without examining the ethical implications or the appropriateness of the source. While it is typically easier to use religious texts, song lyrics, or movie subtitles, these should be carefully considered Hutchinson (2024); Mager et al. (2023). For instance, song lyrics are not representative of daily communication Mayer et al. (2008) since, e.g., they often rhyme and contain profanities. Additionally, using religious texts without a thorough inspection of potential implications can lead to misrepresentations (e.g., considering some cultures to be highly religious) Mager et al. (2023). Further, some work turns to synthetic data generated using machine translation and LLMs when these show clear limitations, especially in multicultural settings Hershcovich et al. (2022)(see Section 4.2.1). Research from other disciplines, even tightly related, such as linguistics Turner (2023), can help us choose adequate and suitable data sources.
6 Conclusion
We present insights from NLP researchers and practitioners working on under-served languages. We discuss limitations and research practices in the field, then provide actionable recommendations on how to address the reported issues while remaining fair to data workers. Our work is the first to document NLP researchers and workers’ experiences.
Limitations
We acknowledge the fact that there are experiences that are different from those of our respondents and the risk of selection bias. Nonetheless, it is also important to give voice to the concerns of data annotators and researchers working on mid- to low-resource languages, and our survey and this paper aim to do that.
Ethical Considerations
While most respondents shared their contact information, it was mainly for following up on the resulting study. We do not share any information that may reveal their identity or the projects they reported.
7 Acknowledgments
We thank our survey respondents for providing insightful comments and we are grateful to anyone who helped us spread the word on social media or by email.
Many thanks to Joanne Boisson, Michael Schlichtkrull, Sowmya Vajjala, and Evani Radiya-Dixit for their helpful feedback.
Appendix A Appendix
A.1 Questionnaire
We would like to investigate the common practices in NLP research on low-resource languages (language variants and "dialects" included).
If you are/were involved in NLP research on low-resource languages, we would like to hear from you. Note that we **will not** share your name or demographic information in public. We will only be checking your name for potential follow-up.
(You can also include your initials if you do not want to disclose your name.)
  • Email.
  • Name.
  • (Optional) Occupation/Affiliation (if any).
  • Which languages do you work on? Language variants and "dialects" included. Please use commas to separate the languages. E.g., language 1, language 2, …
  • What kind of NLP tasks are you interested in? You can name more than one.
  • What kind of NLP tools would be relevant/useful for your language(s)?
  • Why do you work on this/these language(s) ? You can choose more than one option.
    • I have a genuine interest in languages.
    • I want to build technologies for as many languages as possible.
    • I want to build technologies for my language.
    • Existing technologies in my language of interest suffered from marked limitations.
    • I want to contribute to research on LLMs.
    • I have a genuine interest in NLP/CL/ML.
    • Other. [Note that this is a free text field]
  • (Optional) If your answer to the previous question included "Existing technologies in my language of interest suffered from marked limitations.", can you tell us why? You can choose more than one option.
    • Resources are scarce.
    • The data is not representative of the language usage.
    • The annotation is not performed by fluent speakers.
    • The tools do not perform well.
    • The tools are not aligned with the needs of the language speakers.
    • The tools are not that useful.
    • Other. [Note that this is a free text field]
  • (Optional) If you answered "Existing technologies my language of interest suffered from marked limitations.", can you give an example of these resources or tools?
  • (Optional) If you answered "Existing technologies my language of interest suffered from marked limitations.", can you share why?
  • If you were involved in previous projects, what kind of work were you involved in?
    • Annotation.
    • Data collection.
    • Data creation (e.g., coming up with instructions, questions, etc)
    • Other. [Note that this is a free text field]
  • If you were involved in previous projects, did you often get credit for it?
    • Always.
    • Often.
    • Sometimes.
    • Rarely.
    • Never.
    • Other. [Note that this is a free text field]
  • (Optional) If you were involved in the data collection and/or data annotation in previous projects, how often were you offered authorship?
    • Always.
    • Often.
    • Sometimes.
    • Rarely.
    • Never.
    • Other. [Note that this is a free text field]
  • (Optional) In projects for which you did not receive financial compensation or authorship, and where you were involved in the data collection and/or data annotation, what was your incentive?
    • I was part of a community.
    • I had access to additional resources (e.g., GPUs, data, etc.).
    • I was acknowledged on the project website.
    • I was acknowledged in the paper.
    • Other. [Note that this is a free text field]
  • (Optional) For projects where you were simply acknowledged for being an annotator, how long did the data annotation process take?
    • <=2 hours.
    • 2-6 hours.
    • A day of work.
    • 1-7 days.
    • Other. [Note that this is a free text field]
  • Are you part of a community? (Yes/No)
  • (Optional) If you are part of a community, can you name it?
  • (Optional) Were you involved in projects with industry or academia?
    • Industry.
    • Academia.
    • Both.
  • (Optional long text answer) Can you name the institutions/projects? (We will not make the names public if you do not want to share the names publicly. See question below.)
  • Are you happy making the project names public? (Yes/No)
  • (Optional long text answer) What are the potential challenges that the NLP/CL community working on the languages that you mentioned face?
  • Would you like to receive updates about this project? (Yes/No)
A.2 Languages
The full list of the languages that our respondents have worked is included in the following. Note that participants could work on more than one language. Some may have also conducted research for high-resource languages and the respondents may or may not speak the language(s).
Named Mid- to Low-resource Languages
Afaan Oromo, Albanian, Algerian Arabic, Amharic, Assamese, Awigna, Azerbaijani, Bangla, Basque, Bikol, Cebuano, Coptic, Creole, Croatian, Danish, Egyptian Arabic, Emakhuwa, Faroese, Filipino, Geez, Greek, Harari, Hausa, Hindi, Igbo, Ilocano, Indonesian, Irish, IsiXhosa, Kanuri, Kazakh, Kinyarwanda, Kiswahili, Korean, Light Warlpiri, Lingala, Luganda, Luhya (Lumarachi dialect), Malaysian English, Marathi, Moroccan Arabic, Nepalese, Nyanja, Oromo, Persian/Farsi, Pidgin, Punjabi, Raramuri Russian, Saudi Arabic, Sena, Setswana, Sundanese, Swahili, Tagalog, Tarifit Berber, Tigrinya, Tsonga, Tunisian Arabic, Turkish, Urdu, Warlpiri, Welsh, Wixarika, Wolof, Xhosa, Yoreme Nokki, Yorùbá, Zulu.
Families of Languages
African languages, Arabic dialects/variations, English variants, Chatino languages, Gaeilge (including all dialects), Latin American Spanish, Indian languages, Indonesian languages, Nahuatl languages, North African dialects, South East Asian languages.
Named High-resource Languages
English, French, Italian, Modern Standard Arabic, Spanish.